Circular No. 7339

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999gq IN NGC 4523 K. Ayani, Bisei Astronomical Observatory (BAO); and H. Yamaoka,

Kyushu University, report: “We obtained a wide-band spectrum (range 400–800 nm; resolution 0.6 nm) of SN 1999gq (cf. IAUC 7337) with the BAO 1.01-m reflector on Dec. 26.84 UT. This shows that SN 1999gq is of type-II in the early phase, with a broad P-Cyg Hα profile and Hβ absorption on a blue continuum. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity for NGC 4523 (260 km/s), the expansion velocity of the supernova, derived from the Hα absorption minimum, is 7800 km/s. Other broad absorption features are seen at 580, 559, 510, and 495 nm in the rest frame of the host galaxy.”

S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum of SN 1999gq, obtained by P. Berlind on Dec. 29.54 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova after maximum light. The spectrum exhibits strong Balmer features, as well as broad Fe II and Na I lines. Given the NED recession velocity of 262 km/s for NGC 4523, the Hα P-Cyg profile yields an expansion velocity of 7800 km/s for the supernova.

M. Armstrong, Rolvenden, Kent, England, reports that he made an independent discovery of SN 1999gq on a single unfiltered CCD image taken on Dec. 14.23 UT, at mag 14.2. Poor weather prevented him from obtaining a confirming image before the LOSS discovery (cf. IAUC 7337).

SUPERNOVA 1999go IN NGC 1376 M. Salvo, L. Rizzi, A. Pastorello, and M. Turatto, Padova Astronom-

ical Observatory, write: “Spectra (range 370–910 nm; resolution 1.1 nm) of SN 1999go (cf. IAUC 7337) have been taken under poor conditions with the 1.54-m Danish telescope (+ DFOSC) at the European Southern Observatory, La Silla, on Dec. 28.13 and 29.16 UT. The spectra are very blue and featureless on both dates. The slope of the continuum on the second date seems redder, and a hint of a broad absorption centered at 477 nm is visible, probably due to Hβ. SN 1999go is likely a type-II supernova at a very early epoch.”

K. Krisciunas, University of Washington; and K. Loomis, Apache Point Observatory (APO), report preliminary infrared J (1.25-μm) and K′ (2.12μm) photometry for SN 1999go, using the APO 3.5-m telescope (+ GRIM II camera) with a 7′′.6-diameter software aperture: Dec. 27.04 UT, J = 15.25 ± 0.04, K′ = 14.80 ± 0.06.

1999 December 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green